OpenStudy (usukidoll):

Can someone give me a step by step tutorial on substitutions on multiple integrals and using the Jacobian?

OpenStudy (usukidoll):

Question: Use the transformation from exercise 1 to evaluate the integral  for the region R in the first quadrant bounded by the lines y=-2x+4 y =-2x+7 y=x-2 and y=x+1 From exercise 1 solve the system u = x-y v = 2x+y for x and y in terms of u and y. Then find the value of the jacobian.